QUESTION: SOLVE EACH OF THE FOLLOWING SYSTEMS BY SUBSTITUION
20. 8X-4Y=16 I BELIVE YOUR SUBSTITUING FOR Y
Y=2X-4
28. 4X-12Y=5
-X+ 3Y=-1

Answer by stanbon(57290)

20. 8X-4Y=16
Y=2X-4
8x-4(2x-4)=16
8x-8x+16=16
0=0
This means the original problem has an infinite number
of solutions ; The system of equations is true for all
points on the line y=2x-4; it also means you don't have two equations,
you have one.
-------------
28. 4X-12Y=5
-X+ 3Y=-1
x=3Y+1
4(3y+1)-12y=5
12y+4-12y =5
4=5
This is false, of couse. It means there is
no solution to the system of equations. The
two lines are parallel; they never meet in a point.
cheers,
Stan H.